OpenStudy (anonymous):

Need help with this derivative problem! d/dx ln(4x-x^2)

OpenStudy (anonymous):

\[\frac{d}{dx}[\ln(f(x)]=\frac{f'(x)}{f(x)}\]

OpenStudy (anonymous):

wait so I take the derivative and put it over the original problem?

OpenStudy (anonymous):

oh wait nevermind I think I get it. Haha. Just one minute I am going to try and solve it.

OpenStudy (anonymous):

4-2x/4x-x^2????

OpenStudy (anonymous):

@satellite73 ???????

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

Awesome! I just didn't know the formula. :)

OpenStudy (anonymous):

Derivative of In(x) is 1/x

OpenStudy (anonymous):

try the chain rule. We know that the derivative of ln(x) is 1/x so we should write 1/(4x-x^2) then we will need to multiply it by the derivative of the inside function (the 4x-x^2) so do that. If you put it all together you will see how it resembles satellite's formula!

OpenStudy (anonymous):

Is the formula the same for this one too?? ln(x(sqrt(x+1)???

OpenStudy (anonymous):

well the derivative of that problem.

OpenStudy (anonymous):

Yes except there will be 3 steps for the chain rule. First take care of the ln, then the square root (which you can change to a fraction exponent) and then the inside function.

OpenStudy (anonymous):

I tried doing the same thing but I got 3x+2/2(sqrt(x+1)/x(sqrt(x+1) and I do not believe that is the correct answer.

OpenStudy (anonymous):

You just need to remember that the derivative of ln(x) = 1/x

OpenStudy (anonymous):

I will try it out

OpenStudy (anonymous):

Whoops I gave you some wrong info. You will actually need to use the product rule on the inside of that function!

OpenStudy (anonymous):

Yeah I am not quite sure how you got to your answer

OpenStudy (anonymous):

Do you want to go through it step by step?

OpenStudy (anonymous):

haha. I am not sure how to get to the right answer. Yes that would be so helpful!

OpenStudy (anonymous):

Okay so first we see there is an ln on the outside, so lets take the derivative of that to get 1/(xsqrt(x+1))

OpenStudy (anonymous):

The answer is 1/x + 1/2(sqrt(x+1)^2

OpenStudy (anonymous):

Could someone please verify if it is right?

OpenStudy (anonymous):

Now we want to take the derivative of the stuff that is INSIDE the ln function. We can see that we will need to use the product rule. Have you familiar with it?

OpenStudy (anonymous):

inside okay let me see if I can solve it....

OpenStudy (anonymous):

the (x+1) part?

OpenStudy (anonymous):

think of the square root of (x+1) as (x+1)^(1/2)

OpenStudy (anonymous):

so we have x(x+1)^(1/2) and we want to find the derivative of these two things

OpenStudy (anonymous):

we can't exactly multiply that x into the (x+1)^(1/2), so we must use the product rule which is d/dx f(x)g(x)=f(x)'g(x)+f(x)g(x)'

OpenStudy (anonymous):

I got 3x+2/s(x+1)^1/2

OpenStudy (anonymous):

Is that your full answer?

OpenStudy (anonymous):

3x+2/(2(x+1)^1/2

OpenStudy (anonymous):

no just for that inside part

OpenStudy (anonymous):

ok I will try it

OpenStudy (anonymous):

okay I think you have made a mistake somewhere in there. Did you use the product rule?

OpenStudy (anonymous):

I don't know what rule I used. I used the one that I know.

OpenStudy (anonymous):

the f(x)g(x)=f(x)'g(x)+f(x)g(x)' one?

OpenStudy (anonymous):

that is meant to say the derivative of f(x)g(x)

OpenStudy (anonymous):

Does that look familiar? You make know it as 'take the derivative of the first times the second... plus the derivative of the second times the first'

OpenStudy (anonymous):

oh ya I know that one

OpenStudy (anonymous):

and do you know how to take the derivative when there are exponents involved?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

when taking the derivative of the (x+1)^(1/2) move the exponent (1/2) over to the front of the (x+1) and subtract 1 from the exponent. This should probably seem really familar to you. The only difference is that now we have more stuff under the exponent (we have an x+1) so we must multiply what we get by the DERIVATIVE of the x+1

OpenStudy (anonymous):